On two investments totaling $8,000, Lydia lost 4% on one and earned 6% on the other. If her net annual receipts were $400, how much was each investment?
Answer by richwmiller(17219)

Total amount of money invested: $8000
x+y=8000,
Total yearly interest for the two accounts is: $400
-0.04*x+0.06*y=400
x=8000-y
Substitute for x
-0.04*(8000-y)+0.06*y=400
Multiply out
-320--0.04y+0.06*y=400
Combine like terms.
0.1*y=720
Isolate y
y=$ 7200.00 at 6%
x=8000-y
Calculate x
x=$ 800.00 at -4% at a loss
Check
-0.04*800+0.06*7200=400
-32+432=400
400=400
If this statement is TRUE and neither x nor y is negative then all is well